INAF queue observing, 2017 Apr 15-16

Observer: R. Carini, M. Faccini
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: MODS; LUCI

Summary

At the sunset the sky is clear. We can not use LBC, there are problems in the blue camera!!!!

Night duration (12\xB0 el twilight) 03:00am-11:50pm UT = 08 h 50m

Observing time: 08 hrs 5 m (91.5%)

Weather time loss: 0 hrs (0%)

technical time loss 45 m (8.5%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Grazian C cosmos 775 dual grating 2 h 30 m completed
SpecPhot feige 34 dual grating    
Zappacosta A J1521 dual grating 400 s completed
telluric Ricci A3 HIP56736 J    
Ricci A3 SDSSJ11 J 1h 30m completed
telluric A4 HIP61471 J    
Ricci A4 NGC4258 J 1 h 30 m with LUCI1 maybe to be repeated
telluric Piconcelli O HIP68868 HKspec    
Piconcelli O 1451+1441 HKspec 42 m completed
Standard FS149 Ks    
Moscadelli g074 Ks 10m completed
Sky Flat   Ks    
Flat fild +arc LUCI      

Details

  • Sky Flat field with LBC
  • problem with the blue camera!!! The TO fix the problem
  • good some images in V bessel (red channel)
  • problem again with the blue camera, we can not use LBC!!!!!!!! we switch to MODS

03:00 Grazian program, target C
  • acqBinomods cosmos775_ls.acq mods1r.20170416.0001-0003 , mods2r.20170416.0001-0005

03:29 execBinoMODS cosmos777_ls_3.obs
  • seeing about 0.9''
  • mods1r.0004-0007, mods1b.0001-0004 , mods2r.20170416.0006-0009, mods2b.0001-0004
  • 03:54 seeing about 0.6''0.7''
  • 04:28 seeing about 0.7''-0.8''

04:40 execBinoMODS cosmos777_ls_4.obs
  • mods1r.0008, mods1b.0005 , mods2r.20170416.0010, mods2b.0005
  • seeing about 0.8''
  • presence of thin clouds

05:00 switch to standard star feige34
  • acqBinoMODS feige34.acq mods1r.20170416.0009-0010 , mods2r.20170416.0011-0012

05:14 execBinoMODS feige34.obs mods1r.0011-0013, mods1b.0005-0007 , mods2r.20170416.0013-0015, mods2b.0005-0007

05:30 switch to Zappacosta program
  • acqBinoMODS J1521_ls.acq mods1r.20170416.0014-0016 , mods2r.20170416.0016-0018
  • the clouds are becoming ticker.

05:44 execBinoMODS J1521_ls.obs mods1r.0017-0018, mods1b.0008-0009 , mods2r.20170416.0019-0020, mods2b.0008-0009
  • seeing about 1.1''

05:56 switch to LUCI

06:20 Ricci Program, telluric star for Target A3, HIP56736
  • luci1.2170416.0011-0013 , luci2.20170416.0017-0019 acquisition
  • luci1.2170416.0014-0015 , luci2.20170416.0020-0021 science

06:35 target A3
  • 06:38 error during the commit
  • 06:42 in the script the PA= 389 degree, this error is fixed
  • the guide star is a galaxy, is not good, we change it

07:00 load again the script
  • luci2.20170416.0022-0024 luci1.2170416.0016-0018
  • during the acquisition of the source in LUCI1 the software stuck the
  • 07:14 start again the software
  • we continue the acquisition but we need to manually align luci1
  • 7:26 it's impossible to correct the alignment from software
  • the SA in working in real time to de-freez luci
  • luci2.20170416.0025-0040

07:37 the script is working
  • luci1.2170416.0041-0049 , ==luci2.20170416.0019-0033==science
  • the sky is clearer, just thin clouds

08:31 slew to the telluric for A4 target, HIP61471
  • luci1.2170416.0050-0052 , luci2.20170416.0034-0037 acquisition
  • luci1.2170416.0053-0054 , luci2.20170416.0038-0039 science

08:49 slew to the target A4, NGC4258
  • luci1.2170416.0055-0058 , luci2.20170416.0040-0043
  • 09:00 during the alignment luci freezing again, we close the software (with xkill) and open it again
  • it is frozen again, we close the software and open it again
  • it is stuck again, we change station
  • it is is stuck again
  • we close it and and open it again

09:25 now it works. sometimes the system freezing when the window of the pixel scale is open during the alignment
  • 09:30 go the script
  • luci1.2170416.0059-0064 , luci2.20170416.0043-0049 science
  • 09:58 seeing 0.9''
  • the spectra in LUCI1 maybe are not good

10:15 slew to Piconcelli program, run the telluric star for the target O
  • luci1.2170416.0065-0067 , luci2.20170416.0050-0052 acquisition
  • seeing about 0.7''
  • luci1.2170416.0068-0069 , luci2.20170416.0053-0054 science
  • seeing about 0.8'''

10:32 run the target O
  • luci1.2170416.0070-0072 , luci2.20170416.0055-0078 acquisition
  • luci1.2170416.0073-0079 , luci2.20170416.0079-0084 science

11:24 switch to Moscadelli program, standard FS149 in Ks filter
  • luci1.2170416.0080-0083 , luci2.20170416.0085-0088 science

11:40 telluric HIP71172, for the target of Zappacosta program
  • we do not finish the acquisition because the target is coming out of specific

11:50 target g074 of Moscadelli Ks filter
  • luci1.2170416.0087-0092 , luci2.20170416.0092-0097

12:14 sky flat field with LUCI
  • Flat fields
  • archi luci1.2170416.0123-0124 , luci2.20170416.0134-0135 for Piconcelli e Zappacosta
  • luci1.2170416.0127-0128 , ==luci2.20170416.0138-0139==for Piconcelli e Zappacosta
  • luci1.2170416.0134-0135 , ==luci2.20170416.0145-0146==for Piconcelli e Zappacosta
  • flat filter J
  • luci1.2170416.0137-0141 , luci2.20170416.0148-0155 for Ricci
  • archi luci1.2170416.0142-0143 , luci2.20170416.0158-0159 for Ricci
  • archi luci1.2170416.0175-0176 , luci2.20170416.0186-0187 for ricci
  • luci1.2170416.0144-0145 , luci2.20170416.0160-0161 for Ricci
  • flat H luci1.2170416.0150-0154 , luci2.20170416.0166-0170 for Marziani (flat in luci2 to be repeated, the wavelength is 1.28 and not 1.66)
  • archi H (xenon) luci1.2170416.0177-0178 , luci2.20170416.0188-0189 for Marziani
  • archi H (argon) luci1.2170416.0179-0180 , luci2.20170416.0190-0191 for Marziani
  • flat H2 luci1.2170416.0160-0164 , luci2.20170416.0171-0175 for Moscadelli
  • flat Ks luci1.2170416.0165-0169 , luci2.20170416.0176-0180 for Moscadelli
Topic revision: r9 - 18 Apr 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback